OpenStudy (anonymous):

number for 911??????

OpenStudy (anonymous):

Evaluate 40/12. Show work with steps. (3 points)

OpenStudy (skullpatrol):

That is already a value?

OpenStudy (anonymous):

my paper says the nswer is 69 but idk how

OpenStudy (skullpatrol):

Did you mean "simplify"?

OpenStudy (anonymous):

yea

OpenStudy (skullpatrol):

First factor the numerator and the denominator

OpenStudy (anonymous):

my calc says its 3.333 but its wrong

OpenStudy (skullpatrol):

$$\Huge\frac{40}{12}=\frac{?}{?}$$

OpenStudy (anonymous):

so ? divided by ? = 1 cuz same numb divided by sme number???

OpenStudy (skullpatrol):

$$\Huge\frac{40}{12}=\frac{2*20}{2*6}$$ right?

OpenStudy (skullpatrol):

@helpmeepleaseee.. right?